Vapour density of N2O4 at 60 degree centigrade is found to be 30.6.Find out degree of dissociation of N2O4.


Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

VD=30.6

Molecular weight=2*30.6=61.2

Normal MW=28+64=92

so degree of dissociation=92/61.2=1.5

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com